OpenStudy (dcolley):

What is the value of 2 sin 60°cos30°

OpenStudy (anonymous):

2(\[(\sqrt{3}/2)(\sqrt{3}/2)\]

OpenStudy (mertsj):

\[2\frac{\sqrt{3}}{2}(\frac{\sqrt{3}}{2})=\frac{3}{2}\]

OpenStudy (anonymous):

2(\[\sqrt{9}/2\]

OpenStudy (anonymous):

nvm @Mertsj is correct!!

OpenStudy (anonymous):

the answer is 1.5 if u do it in the calculator which is 3/2 also!!

OpenStudy (anonymous):

My dear friend! the answer to this problem is 3/2

OpenStudy (dcolley):

Thanks you guys!!